# Cohomological rank functions and syzygies of projective bundles on abelian varieties

### v3

@@ 4. Projective Bundles on Abelian Varieties
(4.1) $0=E_{1}\subset E_{2}\subset\cdots\subset E_{\mathfrak{r}}=E,$
such that the quotients $\beta_{t}:=E_{t}/E_{t-1}$ lie in $\operatorname{Pic}^{0}(X)$, for every $t=1,\ldots,\mathfrak{r}$. Let $Y$ be the projective bundle $\mathbb{P}_{X}(E)$ with associated projective morphism $\pi:Y\rightarrow X$. Given a line bundle $A$ with $\mathrm{c}_{1}(A)=\ell$ denote by $\pi_{A}:Y_{A}\rightarrow X$ be the projective bundle associated to the vector bundle $E\times A$. Recall that there is an isomorphism
@@ 4.1. Very ampleness and vanishing properties of $E\otimes A$
By definition the very ampleness of $\mathcal{O}_{Y_{A}}(1)$ is equivalent to the ampleness of the vector bundle $E\otimes A$ on $X$, thus the very ampleness part of our statement is a direct consequence of the following.
@@
Suppose that $X$ and $E$ are as in the statement of the Theorem. Let $\ell$ be a polarization on $X$ satisfying $\beta(\ell)<\frac{1}{p+2}$. Then for every $A$ ample line bundle with $\mathrm{c}_{1}(A)=\ell$ we have that the vector bundles $E_{t}\otimes A$ are $(p+1)$-jet ample. In particular $E\otimes A$ is $(p+1)$-jet ample.
@@ 4.2. Syzygies for tautological bundles
for every $\alpha\in\operatorname{Pic}^{0}(X)$ and every $k\geq 1$, and every non-negative integers $k_{t}$ such that $\sum k_{t}\leq p+1$. Let $j=\sum k_{t}$. By [Chi, Lemma 4.4], we have that
@@ 4.3. Finer results for higher syzygies using the work of Ito
Suppose $(Y,X,\pi)$ as before. Let $A$ be an ample line bundle on $X$ with no base divisor. Let $p\geq 1$. If $\mathcal{I}_{o}\left\langle\frac{r+1}{p+r+2}\ell\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}^{r}$-property. In particular, if $\mathcal{I}_{o}\left\langle\frac{1}{p+2}\ell\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}$ property.
@@ 5. Application to projective bundles on desinglarization of compactified Jacobians
In this section, we investigate syzygies of projective bundles on normalization of compactified Jacobians. Our application provides an improvement of the result presented in [Chi, Theorem 5.4] when it comes to an ample line bundle with no base divisor.
@@
by [BP08b, Proposition 2.1]. So, $\widetilde{J^{d}}(X)$ can be viewed as a $\mathbb{P}^{1}$-bundle over $J^{d}(X_{0})$. For more details about the construction of $\widetilde{J^{d}}(X)$, readers may refer to [BP08b] and [OS79, Proposition 12.1]. Let $E$ be the direct sum $\mathcal{P}_{x}\oplus\mathcal{P}_{z}$ of degree 00 line bundles.

### v4

@@ 4. Projective Bundles on Abelian Varieties
(4.1) $0=E_{0}\subset E_{1}\subset E_{2}\subset\cdots\subset E_{\mathfrak{r}}=E,$
such that the quotients $\beta_{t}:=E_{t}/E_{t-1}$ lie in $\operatorname{Pic}^{0}(X)$, for every $t=1,\ldots,\mathfrak{r}$. Let $Y$ be the projective bundle $\mathbb{P}_{X}(E)$ with associated projective morphism $\pi:Y\rightarrow X$. Given a line bundle $A$ with $\mathrm{c}_{1}(A)=\ell$ denote by $\pi_{A}:Y_{A}\rightarrow X$ be the projective bundle associated to the vector bundle $E\otimes A$. Recall that there is an isomorphism
@@ 4.1. Very ampleness and vanishing properties of $E\otimes A$
By definition the very ampleness of $\mathcal{O}_{Y_{A}}(1)$ is equivalent to the very ampleness of the vector bundle $E\otimes A$ on $X$, thus the very ampleness part of our Theorem A is a direct consequence of the following.
@@
Suppose that $X$ and $E$ are as in the statement of the Theorem A. Let $\ell$ be a polarization on $X$ satisfying $\beta(\ell)<\frac{1}{p+2}$. Then for every $A$ ample line bundle with $\mathrm{c}_{1}(A)=\ell$ we have that the vector bundles $E_{t}\otimes A$ are $(p+1)$-jet ample. In particular $E\otimes A$ is $(p+1)$-jet ample.
@@ 4.2. Syzygies for tautological bundles
for every $\alpha\in\operatorname{Pic}^{0}(X)$ and every $k\geq 1$, and every non-negative integers $k_{t}$ such that $\sum k_{t}\leq p+1$. Let $j=\sum k_{t}$. By [Chi19, Lemma 4.4], we have that
@@ 4.3. Finer results for higher syzygies using the work of Ito
Suppose $(Y,X,\pi)$ as before. Let $A$ be an ample line bundle on $X$. Let $p\geq 1$. If $\mathcal{I}_{o}\left\langle\frac{r+1}{p+r+2}\ell\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}^{r}$-property. In particular, if $\mathcal{I}_{o}\left\langle\frac{1}{p+2}\ell\right\rangle$ is M-regular, then $\mathcal{O}_{Y}(1)\otimes\pi^{*}A$ satisfies $N_{p}$ property.
@@ 5. Application to projective bundles on desinglarization of compactified Jacobians
In this section, we investigate syzygies of projective bundles on normalization of compactified Jacobians. Our application provides an improvement of the result presented in [Chi19, Theorem 5.4] when it comes to an ample line bundle with no base divisor.
@@
by [BP08, Proposition 2.1]. So, $\widetilde{J^{d}}(X)$ can be viewed as a $\mathbb{P}^{1}$-bundle over $J^{d}(X_{0})$. For more details about the construction of $\widetilde{J^{d}}(X)$, readers may refer to [BP08] and [OS79, Proposition 12.1]. Let $E$ be the direct sum $\mathcal{P}_{x}\oplus\mathcal{P}_{z}$ of degree 00 line bundles.
@@ Acknowledgment
Both authors are grateful to the anonymous referee whose comments greatly improved the exposition.
@@
In addition both authors are grateful to A. Ito for his valuable comments to a previous version of this manuscript and for pointing out that our Theorem A was a particular case of a result of his.